OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!???????????????Simplify the expression for all appropriate values of x?? Sq.root 12x+15x^2 DIVIDED BY Sq.root 3x

OpenStudy (hba):

\[\frac{ \sqrt{12x+15x^2} }{ \sqrt{3x} }\]

OpenStudy (anonymous):

CAn't you reduce that?

OpenStudy (hba):

Yes you can Try taking common from the numerator :)

OpenStudy (anonymous):

\[\sqrt{4+5x}\] ???

OpenStudy (hba):

Are you getting that as your answer ?

OpenStudy (hba):

@kaykim95

OpenStudy (anonymous):

yes

OpenStudy (hba):

Looks good to me :)